OpenStudy (anonymous):

some help me to solve this: 1/y+1/y+4=1/7

OpenStudy (anonymous):

\[\frac{1}{y}+\frac{1}{y+4}=\frac{1}{7}\]?

OpenStudy (anonymous):

a good first step would be to add the fractions on the left to get \[\frac{y+4+y}{y(y+4)}=\frac{2y+4}{y(y+4)}=\frac{1}{7}\]

OpenStudy (anonymous):

then solve the ratio via \[7(2y+4)=y(y+4)\]

OpenStudy (anonymous):

you good from there?

OpenStudy (anonymous):

I believe so ..

OpenStudy (anonymous):

ok you will have to solve a quadratic equation

OpenStudy (anonymous):

so far I have this y^2-10y-28